OpenStudy (anonymous):

Find the integral of e^3x(e^3x +5)^4 dx

OpenStudy (anonymous):

try \(u=e^{3x}+5\)

OpenStudy (anonymous):

then \(\frac{1}{3}du=e^{3x}dx\) and you get \[\frac{1}{3}\int u^4du\]